OpenStudy (imamermaid29):

NEED HELP ASAP The system of equations is coincident. What are the missing values?

OpenStudy (imamermaid29):

3x + y = 7 9x + ___ y = ___

OpenStudy (caozeyuan):

coincident means the same

OpenStudy (imamermaid29):

i know :)

OpenStudy (caozeyuan):

ay=3y,b=3*7=21

OpenStudy (caozeyuan):

where a is the first black, b is the second

OpenStudy (imamermaid29):

so what would go in the blanks?

OpenStudy (caozeyuan):

3 in the first, 21 in the second

OpenStudy (imamermaid29):

thank you a lot!!! i suck at math XD

OpenStudy (imamermaid29):

and i gave you a fan and medal :)